Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
amagazzu - 16:00 Tuesday 25 January 2022 (54598) Print this report
Operator Report - Morning shift

The shift was dedicated to the planned Maintenance. Here is a list of the activities reported to the Control Room:

  • Cleaning of Central Building (Menzione, Ciardelli with external firm);
  • Standard Vacuum Refill (VAC Team);
  • DAQ SDB2_B1p_QD{1,2} demodulated at 50MHz (see entry #54597);
  • From 8:15 UTC to 9:15 UTC, CERN tiltmeter maintenance at NE Building (Di Girolamo);
  • At around 8:30 UTC, replacement of Nexus power supply in TCS Room (Tringali);
  • From 8:30 UTC to 11:50 UTC, ENV Intervention at NE Building (Fiori, Tringali, Paoletti);
  • From 11:15 UTC to 11:50 UTC, TCS Inspection at NE (Lumaca, Nardecchia);
  • At 9:25 UTC, TCS chiller refill (Menzione);
  • DAQ trend stream (see entry #54599).

From 14:15 UTC Boschi worked to restore the missing connection with Sa_NE (see Network and SUSP sections), activity still in progress.

Sub-system reports

At 11:24 UTC we lost the connection with the imms-13 box (see "Network" section). After the switch was restored I unplugged and plugged back the Ethernat cable of the device: this action restore its connection and the ENV signals were acquired again.

Network & GSS
At 11:24 UTC signals related to the NE started to be corrupted (ENV signals related to Hall temperature and lights) or missing (Sa_NE Loop and Guardians).
At 11:50 UTC I went to investigate the issue at the building and found that one of the Network Switch was off: after pressing its electrical outlet lightly in the socket it turned on again. After this action the connected devices restored their network connections. Expert informed via phone of these events.

SNEB SBE opened multiple times during the shift due to human presence in the Building. Loop closed from VPM.

Due to the issue at NE (see "Network" section) the signals related to NE_IP Loops status and Guardians weren't available anymore from 11:24 UTC. After restoring the Network switch the signals were still missing, SUS_NE Metatron was in Error state and unable to change state. From 14:15 UTC Boschi worked to restore the connection of Sa_NE, activity still in progress.

Images attached to this report
Search Help